Clayton Industries is planning its operations for next year. Ronnie Clayton, the CEO, wants you to forecast the firm's additional funds needed (AFN). Data for use in your forecast are shown below. Based on the AFN equation, what is the AFN for the coming year? Last year's sales = S0 $350
Last yr's accounts payable $40 Sales growth rate = g 30% Last yr's notes payable $50 Last year's total assets = A0* $360 Last yr's accruals $30 Last year's prof margin = PM 5% Target payout ratio 60%. a. $67.0 b. $78.7 c. $63.9 d. $77.9
1
Expert's answer
2020-05-17T17:45:07-0400

AFN = Projected increase in assets – spontaneous increase in liabilities – any increase in retained earnings

"increase in assets=360\times0.3=108"

"increase in liabilities=(40+30)\times0.3=21"

"increase in retained earnings=350\times1.3\times0.05\times(1-0.6)=350\times1.3\times0.05\times0.4=9.1"

AFN=108-21-9.1=77.9


answer d.